# Mode stability of self-similar wave maps without symmetry in higher dimensions

### v1

@@ B.1. Passing to an equivalent bilinear form
$X=\begin{cases}\begin{pmatrix}A&B\\ C&-A^{t}\end{pmatrix},\quad A\in M_{n}(\mathbb{C}),B,C\in\mathfrak{so}(n)&\text{if }d=2n\\ \begin{pmatrix}A&B&z_{1}\\ C&-A^{t}&z_{2}\\ -z_{2}^{t}&-z_{1}^{t}&0\end{pmatrix},\quad A\in M_{n}(\mathbb{C}),B,C\in\mathfrak{so}(n),z_{1},z_{2}\in\mathbb{C}^{n}&\text{if }d=2n+1\end{cases}.$
Note that we can consider $\mathfrak{so}(\widetilde{b}_{2n})$ as a subspace of $\mathfrak{so}(\widetilde{b}_{2n+1})$.
@@ B.3. Irreducible representations
(B.2) $\mathbb{Y}_{l}^{d}\cong\begin{cases}\Gamma((\ell+1)L_{1})\oplus\Gamma(\ell L_{1})\oplus\Gamma((\ell-1)L_{1}),&d=3\\ \Gamma((\ell+1)L_{1})\oplus\Gamma(\ell L_{1}+L_{2})\oplus\Gamma(\ell L_{1}-L_{2})\oplus\Gamma((\ell-1)L_{1}),&d=4\\ \Gamma((\ell+1)L_{1})\oplus\Gamma(\ell L_{1}+L_{2})\oplus\Gamma((\ell-1)L_{1}),&d\geq 5\end{cases}.$
@@
It remains to be shown that this sum exhausts $\mathbb{Y}_{\ell}^{d}$. This follows since the dimensions of the summands in the decomposition add up to the dimension of $\mathbb{Y}_{\ell}^{d}$, which can be checked using Weyl’s dimension formula, see [20, Corollary 24.6]. ∎
@@
Now we can use this to find the value by which $\sum_{1\leq j<k\leq d}\rho_{\ell}(F_{jk})^{2}$ acts on the irreducible components. One has to be a bit careful with normalization here, since the Killing form $\kappa$ on $\mathfrak{so}(d)$ is a multiple of the trace form

### v2

@@ B.1. Passing to an equivalent bilinear form
$X=\begin{cases}\begin{pmatrix}A&B\\ C&-A^{t}\end{pmatrix},\quad A\in M_{n}(\mathbb{C}),B,C\in\mathfrak{so}(n)&\text{if }d=2n\\ \begin{pmatrix}A&B&z_{1}\\ C&-A^{t}&z_{2}\\ -z_{2}^{t}&-z_{1}^{t}&0\end{pmatrix},\quad A\in M_{n}(\mathbb{C}),B,C\in\mathfrak{so}(n),z_{1},z_{2}\in\mathbb{C}^{n}&\text{if }d=2n+1\end{cases},$
see [20, p. 269]. Note that we can consider $\mathfrak{so}(\widetilde{b}_{2n})$ as a subspace of $\mathfrak{so}(\widetilde{b}_{2n+1})$.
@@ B.3. Irreducible representations
For illustration, we show that $L$ and $\vec{p}$ satisfy the conditions of B.1 for the case $d=3$ and $\ell=1$. First of all, we have $L=a_{1}L_{1}$ for $a_{1}=0,1,2$ which evidently satisfy $a_{1}\geq 0$, hence $L$ is a dominant integral weight. Next, we recall that $\widetilde{\rho}_{1}=\rho_{1}\circ\varphi^{-1}$, where
$\varphi^{-1}(X)=TXT^{-1}=\begin{pmatrix}1&\frac{1}{2}&0\\ i&-\frac{1}{2}i&0\\ 0&0&1\end{pmatrix}X\begin{pmatrix}\frac{1}{2}&-\frac{1}{2}i&0\\ 1&i&0\\ 0&0&1\end{pmatrix}$.
Since $d=3=2\cdot 1+1$, we have $n=1$ and hence $\mathfrak{h}=\text{span}\{H_{1}\}$ where
$H_{1}=\begin{pmatrix}1&0&0\\ 0&-1&0\\ 0&0&0\end{pmatrix}$.
In particular, the only positive root $L_{1}\in\mathfrak{h}^{\ast}$ is given simply by $L_{1}(H_{1})=1$. The root space of this root is spanned by
$X_{1}:=\begin{pmatrix}0&0&1\\ 0&0&0\\ 0&-1&0\end{pmatrix}$.
Now we compute
$\widetilde{\rho}_{1}(H_{1})$ $=i(E_{21}-E_{12})+i(\xi_{2}\partial_{\xi^{1}}-\xi_{1}\partial_{\xi^{2}})$
$\widetilde{\rho}_{1}(X_{1})$ $=E_{13}-E_{31}+i(E_{23}-E_{32})+\xi_{1}\partial_{\xi^{3}}-\xi_{3}\partial_{\xi^{1}}+i(\xi_{2}\partial_{\xi^{3}}-\xi_{3}\partial_{\xi^{2}}),$
where the differential operators act diagonally. Now we just insert $\vec{p}$ in the various cases and compute the relations.
  • $L=2L_{1}$: Here we have $\vec{p}=\begin{pmatrix}\xi_{1}+i\xi_{2}\\ i(\xi_{1}+i\xi_{2})\\ 0\end{pmatrix}$ and we compute
    $\widetilde{\rho}_{1}(H_{1})\vec{p}$ $=\left(\begin{pmatrix}0&-i&0\\ i&0&0\\ 0&0&0\end{pmatrix}+i(\xi_{2}\partial_{\xi^{1}}-\xi_{1}\partial_{\xi^{2}})\right)\begin{pmatrix}\xi_{1}+i\xi_{2}\\ i(\xi_{1}+i\xi_{2})\\ 0\end{pmatrix}$
    $=\begin{pmatrix}\xi_{1}+i\xi_{2}\\ i(\xi_{1}+i\xi_{2})\\ 0\end{pmatrix}+\begin{pmatrix}i(\xi_{2}-i\xi_{1})\\ -(\xi_{2}-i\xi_{1})\\ 0\end{pmatrix}=\begin{pmatrix}\xi_{1}+i\xi_{2}\\ i(\xi_{1}+i\xi_{2})\\ 0\end{pmatrix}+\begin{pmatrix}\xi_{1}+i\xi_{2}\\ i(\xi_{1}+i\xi_{2})\\ 0\end{pmatrix}=2\vec{p}=L(H_{1})\vec{p}$
    $\widetilde{\rho}_{1}(X_{1})\vec{p}$ $=\left(\begin{pmatrix}0&0&1\\ 0&0&i\\ -1&-i&0\end{pmatrix}+\xi_{1}\partial_{\xi^{3}}-\xi_{3}\partial_{\xi^{1}}+i(\xi_{2}\partial_{\xi^{3}}-\xi_{3}\partial_{\xi^{2}})\right)\begin{pmatrix}\xi_{1}+i\xi_{2}\\ i(\xi_{1}+i\xi_{2})\\ 0\end{pmatrix}$
    $=\begin{pmatrix}0\\ 0\\ -(\xi_{1}+i\xi_{2})+\xi_{1}+i\xi_{2}\end{pmatrix}+\begin{pmatrix}-\xi_{3}\\ -i\xi_{3}\\ 0\end{pmatrix}+\begin{pmatrix}i(-i\xi_{3})\\ -(-i\xi_{3})\\ 0\end{pmatrix}$
    $=\begin{pmatrix}0\\ 0\\ 0\end{pmatrix}+\begin{pmatrix}-\xi_{3}\\ -i\xi_{3}\\ 0\end{pmatrix}+\begin{pmatrix}\xi_{3}\\ i\xi_{3}\\ 0\end{pmatrix}=0.$
  • $L=L_{1}$: Here we have $\vec{p}=\begin{pmatrix}-\xi_{3}\\ -i\xi_{3}\\ \xi_{1}+i\xi_{2}\end{pmatrix}$ and we compute
    $\widetilde{\rho}_{1}(H_{1})\vec{p}$ $=\left(\begin{pmatrix}0&-i&0\\ i&0&0\\ 0&0&0\end{pmatrix}+i(\xi_{2}\partial_{\xi^{1}}-\xi_{1}\partial_{\xi^{2}})\right)\begin{pmatrix}-\xi_{3}\\ -i\xi_{3}\\ \xi_{1}+i\xi_{2}\end{pmatrix}$
    $=\begin{pmatrix}-\xi_{3}\\ -i\xi_{3}\\ 0\end{pmatrix}+\begin{pmatrix}0\\ 0\\ i(\xi_{2}-i\xi_{1})\end{pmatrix}=\begin{pmatrix}-\xi_{3}\\ -i\xi_{3}\\ 0\end{pmatrix}+\begin{pmatrix}0\\ 0\\ \xi_{1}+i\xi_{2}\end{pmatrix}=\vec{p}=L(H_{1})\vec{p}$
    $\widetilde{\rho}_{1}(X_{1})\vec{p}$ $=\left(\begin{pmatrix}0&0&1\\ 0&0&i\\ -1&-i&0\end{pmatrix}+\xi_{1}\partial_{\xi^{3}}-\xi_{3}\partial_{\xi^{1}}+i(\xi_{2}\partial_{\xi^{3}}-\xi_{3}\partial_{\xi^{2}})\right)\begin{pmatrix}-\xi_{3}\\ -i\xi_{3}\\ \xi_{1}+i\xi_{2}\end{pmatrix}$
    $=\begin{pmatrix}\xi_{1}+i\xi_{2}\\ i\xi_{1}-\xi_{2}\\ \xi_{3}-\xi_{3}\end{pmatrix}+\begin{pmatrix}-\xi_{1}\\ -i\xi_{1}\\ -\xi_{3}\end{pmatrix}+\begin{pmatrix}-i\xi_{2}\\ \xi_{2}\\ \xi_{3}\end{pmatrix}=0.$
  • $L=0L_{1}$: We have $\vec{p}=2i\xi=\begin{pmatrix}2i\xi_{1}\\ 2i\xi_{2}\\ 2i\xi_{3}\end{pmatrix}$ and we compute
    $\widetilde{\rho}_{1}(H_{1})\vec{p}$ $=\left(\begin{pmatrix}0&-i&0\\ i&0&0\\ 0&0&0\end{pmatrix}+i(\xi_{2}\partial_{\xi^{1}}-\xi_{1}\partial_{\xi^{2}})\right)\begin{pmatrix}2i\xi_{1}\\ 2i\xi_{2}\\ 2i\xi_{3}\end{pmatrix}$
    $=\begin{pmatrix}2\xi_{2}\\ -2\xi_{1}\\ 0\end{pmatrix}+\begin{pmatrix}i(2i\xi_{2})\\ i(-2i\xi_{1})\\ 0\end{pmatrix}=\begin{pmatrix}2\xi_{2}\\ -2\xi_{1}\\ 0\end{pmatrix}+\begin{pmatrix}-2\xi_{2}\\ 2\xi_{1}\\ 0\end{pmatrix}=0=L(H_{1})\vec{p}$
    $\widetilde{\rho}_{1}(X_{1})\vec{p}$ $=\left(\begin{pmatrix}0&0&1\\ 0&0&i\\ -1&-i&0\end{pmatrix}+\xi_{1}\partial_{\xi^{3}}-\xi_{3}\partial_{\xi^{1}}+i(\xi_{2}\partial_{\xi^{3}}-\xi_{3}\partial_{\xi^{2}})\right)\begin{pmatrix}2i\xi_{1}\\ 2i\xi_{2}\\ 2i\xi_{3}\end{pmatrix}$
    $=\begin{pmatrix}2i\xi_{3}\\ -2\xi_{3}\\ -2i\xi_{1}+2\xi_{2}\end{pmatrix}+\begin{pmatrix}-2i\xi_{3}\\ 0\\ 2i\xi_{1}\end{pmatrix}+\begin{pmatrix}0\\ i(-2i\xi_{3})\\ i(2i\xi_{2})\end{pmatrix}$
    $=\begin{pmatrix}2i\xi_{3}\\ -2\xi_{3}\\ -2i\xi_{1}+2\xi_{2}\end{pmatrix}+\begin{pmatrix}-2i\xi_{3}\\ 0\\ 2i\xi_{1}\end{pmatrix}+\begin{pmatrix}0\\ 2\xi_{3}\\ -2\xi_{2}\end{pmatrix}=0.$
Hence the conditions of B.1 are satisfied in these cases.
@@
(B.2) $\mathbb{Y}_{\ell}^{d}\cong\begin{cases}\Gamma((\ell+1)L_{1})\oplus\Gamma(\ell L_{1})\oplus\Gamma((\ell-1)L_{1}),&d=3\\ \Gamma((\ell+1)L_{1})\oplus\Gamma(\ell L_{1}+L_{2})\oplus\Gamma(\ell L_{1}-L_{2})\oplus\Gamma((\ell-1)L_{1}),&d=4\\ \Gamma((\ell+1)L_{1})\oplus\Gamma(\ell L_{1}+L_{2})\oplus\Gamma((\ell-1)L_{1}),&d\geq 5\end{cases}.$
@@
It remains to be shown that this sum exhausts $\mathbb{Y}_{\ell}^{d}$. This follows since the dimensions of the summands in the decomposition add up to the dimension of $\mathbb{Y}_{\ell}^{d}$, which can be checked using Weyl’s dimension formula, see [20, Corollary 24.6].
We illustrate this for $d=3$ and $\ell=1$. Weyl’s dimension formula reads
$\dim\Gamma(L)=\frac{\prod_{\alpha\in R^{+}}(L+\rho,\alpha)}{\prod_{\alpha\in R^{+}}(\rho,\alpha)},$
where $\rho:=\frac{1}{2}\sum_{\alpha\in R^{+}}\alpha$ is the half-sum of the positive roots and $(\cdot,\cdot)$ is the inner product on $\mathfrak{h}^{\ast}$ induced by the Killing form (see [20, p. 208]). In our case we simply have $R^{+}=\{L_{1}\}$ and hence $\rho=\frac{1}{2}L_{1}$. For $L=a_{1}L_{1}$ we compute
$\dim\Gamma(L)$ $=\frac{\prod_{\alpha\in R^{+}}(L+\rho,\alpha)}{\prod_{\alpha\in R^{+}}(\rho,\alpha)}=\frac{(a_{1}L_{1}+\frac{1}{2}L_{1},L_{1})}{(\frac{1}{2}L_{1},L_{1})}=\frac{\left(a_{1}+\frac{1}{2}\right)}{\frac{1}{2}}\frac{(L_{1},L_{1})}{(L_{1},L_{1})}=2a_{1}+1.$
With this we compute
$\dim\Gamma(2L_{1})+\dim\Gamma(L_{1})+\dim\Gamma(0L_{1})=5+3+1=9=3\cdot 3=\dim(\mathbb{Y}_{1}^{3})$
hence the dimensions coincide as claimed. ∎
@@
We again illustrate this for $d=3$ and $\ell=1$. Eq. $(25.14)$ from [20, p. 418] reads
$c_{\Gamma(L)}=(L+\rho,L+\rho)-(\rho,\rho).$
In our case we have $\rho=\frac{1}{2}L_{1}$ and we compute for $L=a_{1}L_{1}$
$c_{\Gamma(L)}$ $=(L+\rho,L+\rho)-(\rho,\rho)=\left(a_{1}L_{1}+\frac{1}{2}L_{1},a_{1}L_{1}+\frac{1}{2}L_{1}\right)-\left(\frac{1}{2}L_{1},\frac{1}{2}L_{1}\right)$
$=\left(\left(a_{1}+\frac{1}{2}\right)^{2}-\frac{1}{4}\right)(L_{1},L_{1})=a_{1}(a_{1}+1)(L_{1},L_{1}).$
We have the normalization $(L_{1},L_{1})=\frac{1}{2}$ and we compute
$c_{\Gamma(2L_{1})}$ $=\frac{2\cdot 3}{2}=3$
$c_{\Gamma(L_{1})}$ $=\frac{1\cdot 2}{2}=1$
$c_{\Gamma(0L_{1})}$ $=\frac{0\cdot 1}{2}=0,$
which are exactly the values as claimed in (B.3).
Continuing with the general case, we can use this to find the value by which $\sum_{1\leq j<k\leq d}\rho_{\ell}(F_{jk})^{2}$ acts on the irreducible components. One has to be a bit careful with normalization here, since the Killing form $\kappa$ on $\mathfrak{so}(d)$ is a multiple of the trace form